OpenStudy (anonymous):

Which statement is true about the product /2 (3 /2+/18 A. It is rational and equal to 12. B. It is rational and equal to 18. C. It is irrational and equal to 6 + 2 /18 . D. It is irrational and equal to 3 + /18 .

OpenStudy (anonymous):

what is the equation?

OpenStudy (anonymous):

or product? it is not so clear

OpenStudy (anonymous):

sorry lol the slashes are supposed to be square roots :)

OpenStudy (anonymous):

Okay. Let us see then. To do this question we need to know two things. What is irrational and rational, and how do we simplify rational expressions.

OpenStudy (anonymous):

So do you know what irrational and rational numbers in your case are?

OpenStudy (anonymous):

no lol and my mom only gave me 10 min to finish my hw and i still have a lot pls help as fast as you can plssss

OpenStudy (anonymous):

For example, rational number would be \[\sqrt{25} = 5\] An irrational number on the other hand would be something like\[\sqrt{2}\]

OpenStudy (anonymous):

5 more min :/

OpenStudy (anonymous):

So now to our equation. \[\sqrt{2} (3 \sqrt{2}+\sqrt{18})\] We can first simplify the 18. Do you know how?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2 min :O

OpenStudy (anonymous):

10 secs my mom is counting down on me :/

OpenStudy (anonymous):

-_- christ. I am not going to help you by giving you only single answers. So back to our question. To simplify\[\sqrt{18}\] we break it down inside. Remember prime factorization? \[\sqrt{3\times3\times2} = \sqrt{18}\]

OpenStudy (anonymous):

based on that, 3 can be taken out, because it is square root. so now we have \[\sqrt{2}(3\sqrt{2}+3\sqrt{2})\]

OpenStudy (anonymous):

Well she is apparently not interested in learning, so i am moving on.